OpenStudy (dc_girl831):

dc_girl831 You roll a blue number cube and a green number cube. Find P(a number less then 2 on the blue cube and an odd number on the green cube) A.1/4 B.1/6 C.1/3 D.1/12

OpenStudy (faiqraees):

Can you define what type of cube it is?

OpenStudy (sshayer):

1/6*3/6=?

OpenStudy (sshayer):

number<2 is 1 P(blue)=1/6 odd numbers are 1,3,5 P(green)=3/6 reqd probability as solved above.